# The Energy Loss Phenomenon in RLHF: A New Perspective on Mitigating Reward Hacking

### v2

@@ Abstract
This work identifies the Energy Loss Phenomenon in Reinforcement Learning from Human Feedback (RLHF) and its connection to reward hacking. Specifically, energy loss2 in the final layer of a Large Language Model (LLM) gradually increases during the RL process, with an excessive increase in energy loss characterizing reward hacking. Beyond empirical analysis, we further provide a theoretical foundation by proving that, under mild conditions, the increased energy loss reduces the upper bound of contextual relevance in LLMs, which is a critical aspect of reward hacking as the reduced contextual relevance typically indicates overfitting to reward model-favored patterns in RL. To address this issue, we propose an Energy loss-aware PPO algorithm (EPPO) which penalizes the increase in energy loss in the LLM’s final layer during reward calculation to prevent excessive energy loss, thereby mitigating reward hacking. We theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, which provides deeper insights into its effectiveness. Extensive experiments across various LLMs and tasks demonstrate the commonality of the energy loss phenomenon, as well as the effectiveness of EPPO in mitigating reward hacking and improving RLHF performance.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning Large Language Models (LLM) with human preferences, enabling helpful, honest, and harmless responses (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). This process involves Supervised Fine-Tuning (SFT) on human-annotated data, followed by Reinforcement Learning (RL) that maximizes rewards from a learned proxy Reward Model (RM) based on human preference rankings. Such approaches have greatly improved LLM alignment with human expectations, driving advanced AI applications (Touvron et al., 2023; Ouyang et al., 2022).
@@
Despite the success, RLHF faces a challenge known as reward hacking, where the policy model optimization under the proxy RM diverges from true human preference (Gao et al., 2023). The issue arises because the RM is an imperfect proxy for human preferences (Miao et al., 2024). LLMs can exploit these imperfections by generating well-formed but less relevant responses that overfit to reward model-favored patterns, such as excessive redundancy or caution, to achieve high RM scores through manipulation (Coste et al., 2024; Chen et al., 2024b). Such overfitting prioritizes adherence to reward model-favored patterns, rather than accurately capturing user intent, which often undermines the contextual relevance of responses (Zhang et al., 2024). Some typical hacking samples are provided in Appendix K.
@@
In this work, we aim to uncover the underlying mechanisms of reward hacking within LLMs333 Our work focuses on policy models. For clarity, the term “LLMs” in this paper refers specifically to policy models in RLHF. for developing more effective RL regularization techniques. To this end, we investigate the internal representation dynamics of LLMs during the RL process and closely examine the generation of samples prone to hacking. We empirically observe a consistent increase in energy loss in the LLM’s final layer during the RL process, with an excessive increase in energy loss characterizing reward hacking. We formally define the energy loss phenomenon in Definition 2 and provide a visual illustration in Figure 1. This observation aligns with recent advancements in LLM representation engineering, highlighting the critical role of the LLM’s final layer in encapsulating essential information (Zhang et al., 2024).
@@
We demonstrate the widespread presence of the energy loss phenomenon on four popular LLMs (Llama3-8B (Dubey et al., 2024), Llama2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Deepspeek-7B (Bi et al., 2024)) and two representative tasks (general dialogue (Bai et al., 2022) and summarization (Stiennon et al., 2020)). To address this phenomenon, we propose an Energy loss-aware PPO algorithm (EPPO) that mitigates reward hacking by penalizing the increase in energy loss in the LLM’s final layer during reward calculation. Unlike existing RLHF algorithms that mitigate reward hacking by imposing direct constraints on output spaces, EPPO takes a distinct approach by concentrating solely on the energy loss in the final layer. This design enables a broader scope for policy exploration and optimization, leading to superior performance, as shown in Figure 2. Additionally, in Section 6.1, we theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, providing deeper insights into its effectiveness. Our main contributions are as follows:
@@ 2.2 Reward Hacking in RLHF
Given the challenges of achieving robust reward modeling in real-world scenarios (Casper et al., 2023; Wang et al., 2024), designing RL regularizations to mitigate reward hacking is crucial (Ouyang et al., 2022). Existing regularizations often target output space, such as KL divergence and response length penalties (Singhal et al., 2024; Ouyang et al., 2022), failing to address the internal mechanisms of reward hacking in LLMs. This inevitably limits the optimization landscape and often degrades RLHF performance (Azar et al., 2024; Chen et al., 2024b). In this paper, we delve into the underlying mechanisms of reward hacking and develop more effective RL regularizations for reward hacking mitigation.
@@ 3 The Energy Loss Phenomenon
In this section, we introduce the energy loss phenomenon, an underlying mechanism of reward hacking, supported by empirical observations of energy loss dynamics in LLMs during RL and their connection to reward hacking. Furthermore, we provide a theoretical foundation for our findings.
@@ 3.1 Empirical Evidence on Energy Loss Phenomenon
In Figure 4, we illustrate the relationship between energy loss and reward hacking. Hacking samples are identified by GPT-4 following Miao et al. (2024), detailed in Section 5.1. As observed, while RLHF models generally exhibit higher energy loss than SFT models, hacking samples from RLHF models show a more excessive increase in energy loss compared to normal ones. This consistent pattern across various LLMs and datasets, as shown in Appendix C, shows the commonality of energy loss phenomenon ❷ in Definition 2.
@@
3.2 Theoretical Analysis on Energy Loss Phenomenon
We establish a theoretical foundation for the energy loss phenomenon by examining the response-context relevance—a critical aspect of reward hacking (Zhang et al., 2024).
Theorem 3.
where $\sigma$ is a positive constant.
Corollary 4.
Remark I: In Section 6.2, we empirically demonstrate the negative correlation between energy loss in the LLM’s final layer and the contextual relevance of responses, validating our theoretical analysis and the energy loss phenomenon.
@@ 5.1 Setup
GPT-4 Evaluation. We evaluate EPPO by comparing the win ratio of RLHF model responses under EPPO against baselines. GPT-4 is adopted for assessment due to its strong alignment with human evaluations (Chen et al., 2023; Zheng et al., 2024). To mitigate positional bias (Wang et al., 2018), each sample pair is evaluated twice with reversed order. We use the GPT-4 prompt with the highest human agreement in AlpacaEval (Li et al., 2023), as detailed in Appendix J.3. To ensure the reliability of our experiments, we also adopt Claude-3.5-Sonnet and human as evaluators in Appendix F.
@@
5.2 Main Results of RLHF Performance
5.3 Main Results of Reward Hacking Mitigation
@@ 7.2 Representation Engineering for Neural Networks
Neural networks, often perceived as chaotic and opaque, have been shown to develop semantically meaningful internal representations. Early work on word embeddings reveals semantic relationships, compositionality (Mikolov et al., 2013), and biases within text corpora (Bolukbasi et al., 2016). Subsequent studies demonstrate that text embeddings could encode dimensions like commonsense morality without explicit instruction (Schramowski et al., 2019). Recent research has increasingly focused on regulating LLM behavior through internal representations to address attributes such as honesty, fairness, and harmlessness (Zou et al., 2023; Park et al., 2024; Azaria & Mitchell, 2023; Ji et al., 2024; Zhu et al., 2024b). Building on this foundation, our paper shows that reward hacking in RLHF manifests within LLM representations and introduces a tailored mitigating strategy.
@@ Appendix A Proof of Theorem and Corollary
In this section, we further analyze the energy loss distribution of responses before and after RLHF, as well as its relationship to reward hacking, across a diverse range of LLMs and datasets. This analysis corresponds to the second characteristic of the energy loss phenomenon defined in Definition 2; see Figures 9, 10, 11, and 12 for related results. The key findings are: ❶ The excessive increase in energy loss consistently corresponds to reward hacking across all LLMs and datasets, emphasizing the widespread presence of the second characteristic of the energy loss phenomenon in Definition 2.Our EPPO effectively mitigates reward hacking in all settings by suppressing the excessive increase in energy loss, aligning with our motivation for developing EPPO and the hacking mitigation analysis presented in Section 5.3. Notably, despite our best efforts, hacking samples in the summarization task are challenging to identify with simple natural language instruction for GPT-4 precisely. Therefore, unlike the general dialogue task, where GPT-4 serves as the hacking annotator, we rely on the recently proposed representation-based detection method, InfoRM, to identify hacking samples in the summarization task. Related human evaluation results are detailed in Section I.
@@ D.1 GPT-4 Identification
We have presented the distribution of normal and hacking responses from RLHF models trained with PPO and our EPPO across a wide range of LLMs and datasets in Figures 9, 10, 11, and 12. As observed, our EPPO consistently demonstrates significant effectiveness in mitigating reward hacking across all settings, consistent with our analysis in Section 5.3

### v3

@@ Abstract
This work identifies the Energy Loss Phenomenon in Reinforcement Learning from Human Feedback (RLHF) and its connection to reward hacking. Specifically, energy loss2 in the final layer of a Large Language Model (LLM) gradually increases during the RL process, with an excessive increase in energy loss characterizing reward hacking. Beyond empirical analysis, we further provide a theoretical foundation by proving that, under mild conditions, the increased energy loss reduces the upper bound of contextual relevance in LLMs, which is a critical aspect of reward hacking as the reduced contextual relevance typically indicates overfitting to reward model-favored patterns in RL. To address this issue, we propose an Energy loss-aware PPO algorithm (EPPO) which penalizes the increase in energy loss in the LLM’s final layer during reward calculation to prevent excessive energy loss, thereby mitigating reward hacking. We theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, which provides deeper insights into its effectiveness. Extensive experiments across various LLMs and tasks demonstrate the commonality of the energy loss phenomenon, as well as the effectiveness of EPPO in mitigating reward hacking and improving RLHF performance. Code will be available at Energy-Loss-Phenomenon.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning Large Language Models (LLM) with human preferences, enabling helpful, honest, and harmless responses (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022; Zhang et al., 2025; Li et al., 2023b). This process involves Supervised Fine-Tuning (SFT) on human-annotated data, followed by Reinforcement Learning (RL) that maximizes rewards from a learned proxy Reward Model (RM) based on human preference rankings. Such approaches have greatly improved LLM alignment with human expectations, driving advanced AI applications (Touvron et al., 2023; Ouyang et al., 2022).
@@
Despite the success, RLHF faces a challenge known as reward hacking, where the policy model optimization under the proxy RM diverges from true human preference (Gao et al., 2023). The issue arises because the RM is an imperfect proxy for human preferences (Miao et al., 2024). LLMs can exploit these imperfections by generating well-formed but less relevant responses that overfit to reward model-favored patterns, such as excessive redundancy or caution, to achieve high RM scores through manipulation (Coste et al., 2024; Chen et al., 2024b). Such overfitting prioritizes adherence to reward model-favored patterns, rather than accurately capturing user intent, which often undermines the contextual relevance of responses (Zhang et al., 2024a). Some typical hacking samples are provided in Appendix K.
@@
In this work, we aim to uncover the underlying mechanisms of reward hacking within LLMs333 Our work focuses on policy models. For clarity, the term “LLMs” in this paper refers specifically to policy models in RLHF. for developing more effective RL regularization techniques. To this end, we investigate the internal representation dynamics of LLMs during the RL process and closely examine the generation of samples prone to hacking. We empirically observe a consistent increase in energy loss in the LLM’s final layer during the RL process, with an excessive increase in energy loss characterizing reward hacking. We formally define the energy loss phenomenon in Definition 2 and provide a visual illustration in Figure 1. This observation aligns with recent advancements in LLM representation engineering, highlighting the critical role of the LLM’s final layer in encapsulating essential information (Zhang et al., 2024a).
@@
We demonstrate the widespread presence of the energy loss phenomenon on four popular LLMs (Llama3-8B (Grattafiori et al., 2024), Llama2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Deepspeek-7B (Bi et al., 2024)) and two representative tasks (general dialogue (Bai et al., 2022) and summarization (Stiennon et al., 2020)). To address this phenomenon, we propose an Energy loss-aware PPO algorithm (EPPO) that mitigates reward hacking by penalizing the increase in energy loss in the LLM’s final layer during reward calculation. Unlike existing RLHF algorithms that mitigate reward hacking by imposing direct constraints on output spaces, EPPO takes a distinct approach by concentrating solely on the energy loss in the final layer. This design enables a broader scope for policy exploration and optimization, leading to superior performance, as shown in Figure 2. Additionally, in Section 6.2, we theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, providing deeper insights into its effectiveness. Our main contributions are as follows:
@@ 2.2 Reward Hacking in RLHF
Given the challenges of achieving robust reward modeling in real-world scenarios (Casper et al., 2023; Wang et al., 2024; Zhang et al., 2024b), designing RL regularizations to mitigate reward hacking is crucial (Ouyang et al., 2022). Existing regularizations often target output space, such as KL divergence and response length penalties (Singhal et al., 2024; Ouyang et al., 2022), failing to address the internal mechanisms of reward hacking in LLMs. This inevitably limits the optimization landscape and often degrades RLHF performance (Azar et al., 2024; Chen et al., 2024b). In this paper, we delve into the underlying mechanisms of reward hacking and develop more effective RL regularizations for reward hacking mitigation.
@@ 3 The Energy Loss Phenomenon
In this section, we introduce the energy loss phenomenon, an underlying mechanism of reward hacking, supported by empirical observations of energy loss dynamics in LLMs during RL and their connection to reward hacking. Furthermore, we also provide a theoretical foundation for our findings in Section 6.1 .
@@ 3.1 Empirical Evidence on Energy Loss Phenomenon
In Figure 4, we illustrate the relationship between energy loss and reward hacking. Hacking samples are identified by GPT-4, as described in Section 5.1. As observed, while RLHF models generally exhibit higher energy loss than SFT models, hacking samples from RLHF models show a more excessive increase in energy loss compared to normal ones. This consistent pattern across various LLMs and datasets, as shown in Appendix C, shows the commonality of energy loss phenomenon ❷ in Definition 2.
Remark I: In Section 6.1, we establish a theoretical foundation for our findings by proving that the excessive increase in energy loss in the final layers of LLMs tends to suppress contextual relevance, ultimately leading to reward hacking.
@@ 5.1 Setup
GPT-4 Evaluation. We evaluate EPPO by comparing the win ratio of RLHF model responses under EPPO against baselines. GPT-4 is adopted for assessment due to its strong alignment with human evaluations (Chen et al., 2023; Zheng et al., 2024). To mitigate positional bias (Wang et al., 2018), each sample pair is evaluated twice with reversed order. We use the GPT-4 prompt with the highest human agreement in AlpacaEval (Li et al., 2023a), as detailed in Appendix J.3. To ensure the reliability of our experiments, we also adopt Claude-3.5-Sonnet and human as evaluators in Appendix F.
@@
5.2 Main Results of RLHF Performance
5.3 Main Results of Reward Hacking Mitigation
@@ 6 Discussion
6.1 Theoretical Analysis on Energy Loss Phenomenon
We establish a theoretical foundation for the energy loss phenomenon by examining the response-context relevance—a critical aspect of reward hacking (Zhang et al., 2024a).
Theorem 3.
where $\sigma$ is a positive constant.
Corollary 4.
Additionally, in Section 6.3, we empirically demonstrate a negative correlation between the energy loss in the LLM’s final layer and the contextual relevance of its responses, thereby validating both our theoretical analysis and the observed energy loss phenomenon.
Remark III: We also empirically observe that the increase in the final-layer energy loss during RL largely stems from a decrease in the energy of the final-layer output hidden state, measured by its $L_{1}$ norm. Accordingly, the emergence of reward hacking behavior is not only characterized by excessive energy loss, but also by a significant decline in the $L_{1}$ norm of the final output hidden state. Building on this observation, we establish a more general theoretical explanation in Theorem 6: the $L_{1}$ norm of the final output hidden state provides an upper bound on the contextual relevance of the response. As a result, as the final-layer energy loss increases during RL, the $L_{1}$ norm of the final output hidden state correspondingly decreases, compressing the contextual relevance of the response and potentially leading to reward hacking. The formal statement and proof of Theorem 6 are provided in Appendix A.
@@ 6.1 Theoretical Analysis on Energy Loss Phenomenon
6.4 Energy Loss Distribution across All LLM Layers
Attentive readers may wonder how energy loss varies across different LLM layers. To address this question, we present the energy loss distribution across all layers in Figure 8. The results reveal that the final layer consistently exhibits the most pronounced and stable pattern across various LLMs and datasets. We hypothesize that, as the layer closest to the output, it captures richer semantic information, making it a more reliable indicator of overfitting or reward hacking. In contrast, earlier layers may show similar trends, but their behaviors are more susceptible to variations in network architecture, training schemes, and data distribution, leading to less consistent and generalizable patterns.
Refer to caption Refer to caption
           LLM: Llama2-7B   &  Dataset: Anthropic-Helpful            LLM: Llama2-7B   &  Dataset: Anthropic-Harmless
Refer to caption Refer to caption
           LLM: Llama3-8B   &  Dataset: Anthropic-Helpful            LLM: Llama3-8B   &  Dataset: Anthropic-Harmless
Refer to caption Refer to caption
           LLM: DeepSeek-7B   &  Dataset: Anthropic-Helpful            LLM: DeepSeek-7B   &  Dataset: Anthropic-Harmless
Figure 8: Energy loss distribution across all LLM layers before and after RLHF. From top to bottom, the LLMs shown are LLaMA2-7B, LLaMA3-8B, and DeepSeek-7B, respectively. From left to right, the dataset used for response generation are Anthropic-Helpful and Anthropic-Harmless, respectively. Observation: After RLHF, final-layer energy loss consistently increases across all LLMs and datasets, with excessive increase often coinciding with the emergence of reward hacking.
@@ 7.2 Representation Engineering for Neural Networks
Neural networks, often perceived as chaotic and opaque, have been shown to develop semantically meaningful internal representations. Early work on word embeddings reveals semantic relationships, compositionality (Mikolov et al., 2013), and biases within text corpora (Bolukbasi et al., 2016). Subsequent studies demonstrate that text embeddings could encode dimensions like commonsense morality without explicit instruction (Schramowski et al., 2019). Recent research has increasingly focused on regulating LLM through internal representations to address attributes such as honesty, fairness, and harmlessness (Zou et al., 2023; Park et al., 2024; Azaria & Mitchell, 2023; Zhu et al., 2024b). Building on this foundation, our paper shows that reward hacking in RLHF manifests within LLM representations and introduces a tailored mitigating strategy.
@@ 8 Conclusion
Acknowledgements
This research is supported by the National Natural Science Foundation of China under Grant 62431020, the Foundation for Innovative Research Groups of Hubei Province under Grant 2024AFA017, the Fundamental Research Funds for the Central Universities under Grant 2042025kf0030.
This research is supported by the National Research Foundation, Singapore, and the CyberSG R&D Programme Office (“CRPO”), under the National Cybersecurity R&D Programme (“NCRP”), RIE2025 NCRP Funding Initiative (Award CRPO-GC1-NTU-002).
@@ Appendix A Proof of Theorem and Corollary
Theorem 6.
Let $X$ and $Y$ be the input and output random variables of an LLM. For any layer (i.e., transformer block) $\ell$, let $H^{\text{in}}_{\ell}$ and $H^{\text{out}}_{\ell}$ be its input and output hidden state random variables. The contextual mutual information $I(X;Y)$ satisfies:
$I(X;Y)\leq\frac{1}{2\sigma^{2}}\mathbb{E}_{p(h_{\ell}^{out})}[\|H_{\ell}^{out}\|^{2}_{1}]+\log\sqrt{2\pi\sigma^{2}},$
where $\sigma$ is a positive constant.
Proof: Given an input $x$ to the LLM, the input and output hidden states of the $\ell$-th layer are denoted as $h_{\ell}^{in}(x)$ and $h_{\ell}^{out}(x)$, respectively. For simplicity, in the following proof, we abbreviate them as $h_{\ell}^{in}$ and $h_{\ell}^{out}$, representing instances of the random variables $H_{\ell}^{in}$ and $H_{\ell}^{out}$, respectively. Additionally, the mapping of the $\ell$-th layer of the LLM is denoted as $f_{\ell}(\cdot)$.
where $\delta(\cdot)$ denotes the Dirac delta function. By introducing a variational distribution $q(h_{\ell}^{out})$, we have:
$\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))-\log q(h_{\ell}^{out})\right]$
The inequality in the above equation is derived from the non-negativity of the KL divergence. Since $\log x\leq x-1$, we have:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))-1-\log q(h_{\ell}^{out})\right]$ (16)
[…]
Substituting Equation (17) into Equation (16), we obtain:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq-\mathbb{E}_{p(h_{\ell}^{out})}\left[\log q(h_{\ell}^{out})\right]$ (18)
To make the variational upper bound in Equation (18) as tight as possible, we need $q(h_{\ell}^{out})$ to closely approximate $p(h_{\ell}^{out})$. Thus, we make the following assumption about the variational distribution $q(h_{\ell}^{out})$:
$q(h_{\ell}^{out})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{1}{2\sigma^{2}}(\|h_{\ell}^{out}\|_{1}-\mathbb{E}[\|H_{\ell}^{out}\|_{1}])^{2}\right).$ (19)
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq-\mathbb{E}_{p(h_{\ell}^{out})}\left[\log q(h_{\ell}^{out})\right]$ (20)
$=\frac{1}{2\sigma^{2}}\mathbb{E}_{p(h_{\ell}^{out})}\left[(\|h_{\ell}^{out}\|_{1}-\mathbb{E}[\|H_{\ell}^{out}\|_{1}])^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$
$=\frac{1}{2\sigma^{2}}\text{Var}(\|H_{\ell}^{out}\|_{1})+\log\sqrt{2\pi\sigma^{2}}$
$\leq\frac{1}{2\sigma^{2}}\mathbb{E}_{p(h_{\ell}^{out})}[\|H_{\ell}^{out}\|^{2}_{1}]+\log\sqrt{2\pi\sigma^{2}}$
Let $X$ and $Y$ denote the input and output of the LLM, respectively. Since $X\rightarrow H_{\ell}^{in}\rightarrow H_{\ell}^{out}\rightarrow Y$ forms a Markov chain, we have
$I(X;Y)\leq I(H_{\ell}^{in};H_{\ell}^{out})\leq\frac{1}{2\sigma^{2}}\mathbb{E}_{p(h_{\ell}^{out})}[\|H_{\ell}^{out}\|^{2}_{1}]+\log\sqrt{2\pi\sigma^{2}}$ (21)
@@
In this section, we further analyze the energy loss distribution of responses before and after RLHF, as well as its relationship to reward hacking, across a diverse range of LLMs and datasets. This analysis corresponds to the second characteristic of the energy loss phenomenon defined in Definition 2; see Figures 10, 11, 12, and 13 for related results. The key findings are: ❶ The excessive increase in energy loss consistently corresponds to reward hacking across all LLMs and datasets, emphasizing the widespread presence of the second characteristic of the energy loss phenomenon in Definition 2.Our EPPO effectively mitigates reward hacking in all settings by suppressing the excessive increase in energy loss, aligning with our motivation for developing EPPO and the hacking mitigation analysis presented in Section 5.3. Notably, despite our best efforts, hacking samples in the summarization task are challenging to identify with simple natural language instruction for GPT-4 precisely. Therefore, unlike the general dialogue task, where GPT-4 serves as the hacking annotator, we rely on the recently proposed representation-based detection method, InfoRM, to identify hacking samples in the summarization task. Related human evaluation results are detailed in Section I.
@@ D.1 GPT-4 Identification
We have presented the distribution of normal and hacking responses from RLHF models trained with PPO and our EPPO across a wide range of LLMs and datasets in Figures 10, 11, 12, and 13. As observed, our EPPO consistently demonstrates significant effectiveness in mitigating reward hacking across all settings, consistent with our analysis in Section 5.3